# Modeling Item-Level Dynamic Variability with Residual Diffusion for Bundle Recommendation

### v2

@@ 3.6. Training Efficiency Study (RQ3)
Table 4. The statistics of one-epoch training time(seconds) for three backbone models and using RDiffBR on them.
Dataset Youshu NetEase
BGCN 1.46 23.5
BGCN+RDiffBR 1.51 24.5
CrossCBR 1.06 8.7
CrossCBR+RDiffBR 1.11 9.1
BundleGT 22.7 426.2
BundleGT+RDiffBR 23.3 434.8
4. RELATED WORK
4.1. Bundle Recommendation
Bundle recommendation aims to recommend a bundle of items related to a specific theme. Early work employed neural network methods for bundle recommendation (Cao et al., 2017; Chen et al., 2019b; He et al., 2019; Liu et al., 2014). Recently, graph neural networks have become the mainstream technology for bundle recommendation due to its strong ability to encode higher-order information. We can categorize these graph bundle recommendation models based on these three types of representation learning strategies (Sun et al., 2024). $\bullet$ Unified learning. The unified representation learning strategy aims to learn representations of users and bundles within a unified framework. There are some existing research studies that utilized this strategy, which integrated users, items and bundles into a unified user-bundle-item graph (Chen et al., 2019a) or hypergraph  (Yu et al., 2022), and propagated neighboring information to get the final representations of users and bundles. $\bullet$ Separate Learning. The main idea of this strategy is to first perform representation learning and preference prediction upon two views individually, and then fuse two view-specific predictions to obtain the final prediction. This works (Chang et al., 2020; Ke et al., 2023b; Tan et al., 2021; Li et al., 2023; Wei et al., 2023) perform representation learning on several views separately and then fuse the specific predictions of views. $\bullet$ Cooperative Learning. This approach is similar to the separate learning strategy in that it learns user and bundle representations from bundle view and item view but the key difference is that, techniques such as contrastive learning (Zhao et al., 2022; Ma et al., 2022, 2024a; Zou et al., 2024), knowledge distillation (Ke et al., 2023a; Ren et al., 2023) are employed cross views to achieve mutual enhancement.
However, most of bundle recommendation models have achieved remarkable effectiveness for predicting the user’s preference for prebuilt bundles. So they lack the ability to adapt to item-level dynamic variability scenario. It is necessary to study our task for real-world application scenarios.
4.2. Diffusion Models for Recommendation
The emerging diffusion model in generative models, compared with GANs and VAEs, possess stronger stability and representation. Diffusion model plays three main roles in recommendation systems (Lin et al., 2024b): $\bullet$ Diffusion for data augmentation. By generating realistic samples to expand the original training data, such as sequence data (Wu et al., 2023; Liu et al., 2023), multi-field features (Kotelnikov et al., 2023; Zhang et al., 2024), user-item interactions (Liu et al., 2024) and side information (Jiang et al., 2024b, a), diffusion models can enhance the performance of recommendation systems from multiple perspectives. $\bullet$ Diffusion for Representation Enhancement. Diffusion models converts the input data into embeddings to address some key challenges of recommendation. Lots of studies have been proposed such as denoising the implicit user feedback (Zhao et al., 2024), out-of-distribution issue (Zhao et al., 2025) and multi-modal learning (Ma et al., 2024b). $\bullet$ Diffusion as Recommendation Model. When using diffusion models as recommendation models, the input at this stage can be structured datas or embeddings from other encoders. The related research can be classified into three categories: collaborative recommendation (Wang et al., 2023; Hou et al., 2024), context-aware recommendation (Li et al., 2024b; Qin et al., 2024), and other applications (Lin et al., 2024a). $\bullet$ Diffusion for Content Presentation. Diffusion models can further enhance user satisfaction by presenting generated content to target users, rather than conducting item recommendations. Existing research works can be roughly classified into three categories: fashion generation (Xu et al., 2024), advertising creative generation (Yang et al., 2024) and general content generation (Wang et al., 2024).
However, simply applying diffusion model on bundle models is insufficient to address the issue of item-level dynamic variability. There are mainly two reasons: the loss targets of diffusion and bundle recommendation are inconsistency and the approximator needs to be designed base on item-level dynamic variability scenario.
5. Conclusion and Future Work
In this work, to address the issue of item-level dynamic variability, we proposed RDiffBR, which is a model-agnostic generative framework that can help a BR model adapt this scenario. Given the item-level recommendation model of the bundle, RDiffBR actively injects Gaussian noise into these embeddings and iteratively removes the noise during the reverse process. To guide the reverse process, we designed a residual app module. In the inference stage, we used this module to generate new item-level embeddings. A large number of experimental results show that RDiffBR enables the backbone BR model to achieve significant performance improvement in scenarios with dynamic changes of B-I. Future improvements of RDiffBR may include: 1) allowing RDiffBR to be directly inserted into the well-trained bundle model and take effect; 2) Using diffusion to model bundle features as distridution rather than fixed features that have limitations in B-I dynamic variability scenario.
Acknowledgements.

### v3

@@ 13     Take gradient descent step on $\nabla_{\theta}(\mathcal{L})$ to optimize $\theta$;
Table 4: The statistics of one-epoch training time (seconds) for three backbone models and using RDiffBR on them.
From the perspective of time complexity, taking lightweight backbone CrossCBR as an example, the time complexity of graph learning is $\mathcal{O}(|E|Kd)$, where $|E|$ is the number of all edges in U-B and U-I graphs, $K$ is the number of propagation layers, $d$ is the embedding size. However, the diffusion process in RDiffBR costs $\mathcal{O}(|B|d)+\mathcal{O}(|B|d)d’n$, where $|B|$ is the number of bundles, $d$’ is the embedding size of MLP in approximator, $n$ is MLP layer number. It shows that $|E|$ is much greater than $|B|$ (e.g., in iFashion $|E|=1679708>>|B|=27694$). Therefore, our RDiffBR only increases training time about $4\%$ on two GTX 4090.
Conclusion and Future Work
In this work, to address the issue of item-level dynamic variability, we proposed RDiffBR, a model-agnostic generative framework that can help a BR model adapt to this scenario. A number of experimental results show that RDiffBR enables backbone BR models to achieve significant performance improvement under dynamic changes of B-I. Future research will consider extending RDiffBR as a plug-and-play module for well-performing pre-trained bundle recommenders and evaluating its applicability to cold-start bundle recommendation.
Acknowledgments
This work is partially supported by NSFC, China (No.62276196), the research grant (RGPIN-2020-07157) from the Natural Science and Engineering Research Council (NSERC) of Canada and York Research Chairs (YRC) program. The authors also gratefully appreciate the anonymous reviewers for their valuable comments and constructive suggestions that greatly helped to improve the quality of the paper.